Mathematics 243 – Algebraic Structures Definitions Quiz 5 – October 13, 2006

Your Name:

A) (5) What is the definition of the inverse image of a subset T ⊂ B under a mapping f : A → B?

B) (5) How many different integers a satisfy a|20? (Include both positive and negative a.)